OpenStudy (anonymous):

Multiply and simplify -3x^2y^2 • x^3 A. 9x5y2 B. –3x5y6 C. –3x5y2 D. –3x6y2

OpenStudy (camerondoherty):

C

OpenStudy (mathmale):

In the interest of clarity, I'm rewriting your -3x^2y^2 • x^3 as\[-3x^2y^2 • x^3\]

OpenStudy (dan815):

Cameron refrain from just giving out the answers please.

OpenStudy (camerondoherty):

ok sorry

OpenStudy (dan815):

no problem :)

OpenStudy (mathmale):

Now, let's re-write tht as \[-3x^2 * x^3 * y^2\]

OpenStudy (mathmale):

What rule governs this multiplication?\[x^a x^b?\]\[x^a*x^b=x ^{( ?? )}\]

OpenStudy (camerondoherty):

Step 1 : Multiply \[3x^2y^2 \times x^3\] x^2 multiplied by x^3 = x(2 + 3) = x^5 Then you get: 0 - 3x^5y^2 Step 2 : Simplify 0 - 3x^5y^2 Final result : -3x^5y^2